40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 11 18:20:59 2018, gautam, Update, Optical Levers, BS/PRM HeNe is dead BSOL_dead.png
    Reply  Sat Dec 15 20:04:03 2018, gautam, Update, IOO, TT1/TT2 stepping 
       Reply  Wed Dec 19 15:15:56 2018, gautam, Update, IOO, TT1/TT2 stepping 
Message ID: 14345     Entry time: Tue Dec 11 18:20:59 2018     Reply to this: 14362
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: BS/PRM HeNe is dead 

I found that the BS/PRM OL SUM channels were reading close to 0. So I went to the optical table, and found that there was no beam from the HeNe. I tried power-cycling the controller, there was no effect. From the trend data, it looks like there was a slow decay over ~400000 seconds (~ 5 days) and then an abrupt shutoff. This is not ideal, because we would have liked to use the Oplevs as a DC alignment reference during the ventnoI plan to use the AS camera to recover some sort of good Michelson alignment, and then if we want to, we can switch out the HeNe.

*How can I export PDF from NDscope?

Attachment 1: BSOL_dead.png  73 kB  | Hide | Hide all
BSOL_dead.png
ELOG V3.1.3-